# Towards Immersive Mixed Reality Street Play

### v1

@@ 2.3. Collocated Mixed Reality
The focus of MR research has increasingly shifted from single-user experiences to multi-user collaborations, particularly within the CSCW community. Billinghurst coined the term ”Collaborative Mixed Reality” to describe this emerging field (Billinghurst and Kato, 2002). Collaborative MR can be broadly categorized into remote collaboration and face-to-face collocated collaboration. Remote collaboration has been extensively studied (Schäfer et al., 2023). These studies have demonstrated MR’s potential to bridge physical distances and enhance communication in distributed teams. On the other hand, face-to-face collocated collaboration in MR (Billinghurst et al., 2002) has been explored in two main areas: interactions between MRHMD wearers (Nilsson et al., 2009) and collaborative work involving both handheld devices and MRHMDs (Zaman et al., 2023). The latter has particularly focused on bystander inclusion, addressing the challenge of integrating non-HMD users into MR experiences (Zhu and Grossman, 2020; Denning et al., 2014). Research on collocated collaboration between HMD wearers has investigated shared workspace designs, collaborative problem-solving, and social dynamics in MR environments (Sereno et al., 2020).
[…]
@@
In recent years, MRHMDs with passthrough functionality (Itoh et al., 2021) have sparked a notable shift from private, controlled settings to more public, spontaneous environments. These devices, including video and optical see-through technologies like Apple Vision Pro, Magic Leap, Microsoft HoloLens, and Meta Orion, enable users to view their physical surroundings while interacting with virtual elements (Radu et al., 2021b). The rise in MRHMD usage in public settings raises social concerns for both wearer and non-wearers: MRHMD usage. The public acceptance tends to decline significantly (Schwind et al., 2018; Eghbali et al., 2019) due stems from three main factors:
@@
MRHMDs grant wearers capabilities—such as data overlays or video recording—that non-wearers cannot see or fully understand. This imbalance can negatively affect social experiences, privacy, and safety (Chung et al., 2023). Denning et al. (Denning et al., 2014) examined bystander privacy in AR, while O’Hagan et al. (O’Hagan et al., 2021) studied bystanders’ privacy concerns and awareness regarding AR glasses. A notable example of these social tensions is Google Glass, which faced criticism for prioritizing user capabilities over broader social impacts. This led to the term “Glasshole,” illustrating bystanders’ perception of antisocial behavior by Google Glass wearers (Due, 2015).
Embodied Interactions.
MRHMDs often rely on embodied interactions (Handosa et al., 2018), such as mid-air gestures and full-body movements. While these can be intuitive for manipulating virtual objects, they can also conflict with established social norms, resulting in social awkwardness (Plakias, 2024). Moreover, these conspicuous gestures may exacerbate the power imbalance between wearers and bystanders, as onlookers may feel excluded or uncertain about the wearer’s actions in mixed reality.
Safety in Public Spaces.
[…]
When MRHMDs are used “on the go” in urban areas or during street play, trust, security, and privacy become even more pressing (Katins et al., 2024; Millard et al., 2024). Interactions and movement in public traffic environments can pose additional risks, especially when MRHMD wearers navigate between physical and virtual elements. This interaction style intensifies the power imbalance and increases the potential for accidents or social friction with bystanders.
@@ 2.5. Research Gap: Immersive Mixed Reality Street Play
[…]
While the concept of IMRSP has been advocated for several years, there remains a significant gap in research regarding collocated bodily play with MRHMDs in public spaces. To our knowledge, no commercially playable IMRSP games currently exist that can be deployed on real streets with ecological validity. Although extensive empirical research has examined the social implications of smartphone-based street play like Pokémon Go (Alha et al., 2019), our literature review reveals no structured understanding or empirical studies of IMRSP. This under-exploration appears to stem from three main reasons:
2.5.1. Inaccessibility of affordable MRHMDs
As of 2024, full-color video or optical see-through MRHMDs—such as Vision Pro, HoloLens, and Meta’s Orion—remain prohibitively expensive, with prices starting at \$3,500 per device. This high cost poses a significant barrier to collocated multiplayer play, which requires multiple MRHMDs. Consequently, research in this area is typically confined to controlled lab settings, limiting the exploration of in-the-wild applications.
2.5.2. Limited technology for spontaneous multiplayer mixed reality collocation
Multiplayer mixed reality collocation requires low-latency spatial pose transmission—less than 20ms per frame—to prevent noticeable lag (Struye et al., 2020). This necessitates either deploying a WiFi router in the wild or utilizing low-latency cellular networks like 5G, which aren’t ubiquitous. Furthermore, routing data through remote servers between players often results in latency that exceeds user tolerance thresholds.
2.5.3. Social acceptance barriers and lack of established social norms
@@ 3. System Design
[…]
To explore the real-world social implications of IMRSP, we employed a research-through-design approach. We created a game probe called ”Multiplayer Omnipresent Fighting Arena” (MOFA) and deployed these games in real-world scenarios. MOFA serves as an open-source MR game framework that enables collocated, multiplayer bodily play through MRHMDs in the street spontaneously with peer-to-peer local wireless networking.
@@ 3.1.1. Easy to understand and learn
The game probe should be easily understood by city dwellers on the streets. Inspired by the Harry Potter fantasy, particularly wizard dueling, we adopted a ‘magic’ theme for the gameplay. This aligns well with the characteristics of augmented reality media—unseen energy invisible to muggles, who are unaware of magic. This theme draws from popular culture that has been deeply rooted in global consciousness for over three decades. It makes the rules and settings intuitively understandable for many without explanation in “in-the-wild” settings. Even when behaviors might seem unusual in public, they’re explainable and comprehensible because people have been familiarized through Harry Potter movies—they recognize these wand dueling interactions. Players can instantly immerse themselves in imaginative wizard and witch roles.
@@ 3.1.2. Physical Props for Social Affordance
All gameplay for MRHMD wearers in MOFA incorporates a physical prop: a wand. We discovered that the wand provides an intuitive affordance for players to understand the hand-waving gesture used to cast spells. This proves effective for three main reasons: (1) It aligns with the stereotypical image of wizardry popularized by the Harry Potter series, enabling people to easily grasp the interaction without tutorials; (2) The physical shape and weight of the wand facilitate a natural “wave down” gesture, which is more intuitive than any bare-handed gesture for casting spells (which typically require tutorials). With wands, the interaction becomes second nature; (3) The wands help passersby readily comprehend the context of the games, even without seeing the augmented reality imaging overlay. Addressing the social acceptance concerns in Section 2.5.3, the wand serves as a strong social affordance for both players and passersby in public spaces.
3.1.3. Collocated Bodily Play without Physical Contact
@@
As envisioned in promotional videos, MRHMD wearers engage in embodied interactions involving full-body movements, transcending traditional screen-touching gameplay. MOFA exemplifies collocated bodily play with MRHMDs, showcasing interactions ranging from single-player to multiplayer scenarios. The game deliberately avoids physical contact for two key reasons: (1) many existing mixed reality systems can’t detect physical contact between players, making tangible interactions like melee attacks impractical in game design; (2) to prioritize player safety, we designed the game with remote attacks only, eliminating the risk of direct physical contact between players. MOFA strategically incorporates a magic shield in front of players to indicate an attacking target, which allows opponents to aim at the shield, providing a clear target without risking physical harm. Unlike traditional fighting games where avatars show damage through fainting, blood spatter, or falling, MOFA uses shield breaks to provide feedback when an opponent successfully hits the target. This design ensures both safety and clear gameplay feedback.
[…]
@@
MOFA is designed for deployment anywhere, transforming any street into a game arena—hence its name, “omnipresent.” It should operate without Wi-Fi or cellular connections. Mixed reality collocation requires low-latency spatial pose transmission (under 20ms per frame) to prevent noticeable lag. To address the research gap in spontaneous collocation technology outlined in Section 2.5.2, we employed an open-source collocation toolkit (Hu et al., 2023). This toolkit utilizes MultipeerConnectivity [^9] https://developer.apple.com/documentation/multipeerconnectivity, visited on October 25, 2024—the technology powering Apple’s AirDrop—to enable local networking. This innovation allows street play to begin anywhere, even in rural areas lacking cellular signals or Wi-Fi access. The technology also features a rapid QR scanning process for synchronizing coordinates across devices, eliminating the need to pre-scan the physical environment and enabling spontaneous player or spectator participation.
@@
Required Number of Players 1 2 3 or more (One player playing Dragon using mobile)
Duration per Game Round 90 seconds 90 seconds Approximately 5 minutes
@@
[…]
3.2. Gameplay Overview
MOFA includes three collocated bodily gameplay probes (see Figure 1 and Table 1):
3.2.1. “The Training” - a single-player gameplay probe
@@ 3.2. Gameplay Overview
“The Training” is a single-player dueling gameplay where players engage in magical battles with an AI-controlled cartoon character (see Figure 1a). Inspired by Harry Potter’s training scenes [^10] https://harrypotter.fandom.com/wiki/Cast-a-Spell_training_room, visited on October 25, 2024, players take on the role of wizards, honing their wand skills to cast spells and attack virtual ghosts in Hogwarts. This game probe focuses on one spell-casting interaction: detecting “waving down” hand gestures with physical wands. It is a single-player MRHMD gameplay in public spaces.
[…]
@@
The user interaction flow of ”The Training” is as follows: First, the player puts on the Apple Watch while holding a wand and launches the game app on both the watch and iPhone. After selecting ”The Training” mode, the player inserts the iPhone into the HoloKit headset. Once wearing the headset and detecting a floor plane, the player taps the Apple Watch to place the virtual AI cartoon avatar opponent on the floor in front of them. At the audio cue ”1, 2, 3, start!” the round begins, and the AI starts casting spells and dodging. The player dodges incoming spells and counterattacks by waving their watch-wearing hand, which triggers spells through motion detection. Players earn one point by completely breaking their opponent’s mixed reality shield. Each round lasts 90 seconds, and the player with the most shield breaks wins.
3.2.2. “The Duel” - a competitive gameplay probe
@@
“The Duel” is a competitive gameplay mode pitting two players against each other in magical combat. It expands on “The Training” mechanics (see Figure 1b), transitioning from player-versus-AI to player-versus-player. Drawing inspiration from Harry Potter’s classic Wizard Dueling [^11] https://harrypotter.fandom.com/wiki/Duelling, visited on October 25, 2024, players use wands to cast spells and fight each other, incorporating dodgeball-like body movements to avoid attacks. This simulates future scenarios where MRHMD users might encounter each other on the street for impromptu duels.
[…]
@@
The user interaction flow of ”The Duel” is similar to ”The Training” with additional connection and collocation steps: First, two human players launch the game app on both their watches and iPhones. The host player selects ”The Duel” mode and displays a QR code on his/her phone screen. The client player scans the host’s QR code with their iPhone. After confirming the connection, both players insert their iPhones into their HoloKit headsets. Once wearing the headsets, the two players face each other and click their Apple Watches to start the game round. At the audio cue ”1, 2, 3, start!” the players cast spells, dodge, and counterattack. We observed that players often make dramatic movements to avoid being hit. Players earn one point by completely breaking their opponent’s mixed reality shield. Each round lasts 90 seconds, and the player with the most shield breaks wins.
[…]
After the round, players can check their Apple Watch to see their winning points, hitting rate, estimated moving distance, and estimated calories burned. We selected 90-second rounds to create a mild exertion game suitable for players with average fitness levels, burning approximately 10 kilocalories per round (see Figure 2(e)). Note that Apple Watch calorie estimates are not accurate enough for scientific cardio research.
3.2.3. ”The Dragon” - a collaboration and puppeteering gameplay probe
@@
“The Dragon,” designed for three or more players and extending from “The Duel,” introduces a new interaction paradigm (see Figure 1c): puppeteering. In this mode, one handheld AR player—dubbed the “puppeteer”—controls a virtual dragon using their mobile phone, interacting with MRHMD users. Meanwhile, players equipped with MRHMDs become dragon-hunting warriors, casting spells to battle the dragon while using body movements to avoid attacks. This gameplay is inspired by Game of Thrones’ famous Dragon Hunting scenes [^12] https://game-of-thrones-winter-is-coming-game.fandom.com/wiki/Dragon_Hunt, visited on October 25, 2024. It simulates a transitional future where MRHMD users coexist with handheld AR users, allowing for collaborative play between different device types when encountering each other on the street.
[…]
@@
The user interaction flow of ”The Dragon” is more complex than the previous two games, as depicted in Figure 2. The host player puppeteers the dragon on their mobile screen (Device 1 in Figure 2). First, the host selects ”The Dragon” mode and displays a QR code on their phone screen. Three human players (Devices 2, 3, 4 in Figure 2) scan the host’s QR code with their iPhones. After confirming the connection, all players insert their iPhones into their HoloKit headsets. Two spectators (Device 5 in Figure 2) can also scan the QR code with their iPhones to join as spectators. Once wearing the headsets, the three players face the designated play area after confirming with the host player and tap their Apple Watches to indicate readiness. The host player then detects the floor plane and spawns the virtual dragon to begin the battle. At the audio cue ”1, 2, 3, start!”, the round begins. Players cast spells to attack the dragon while dodging its attacks. The host controls the dragon using the mobile screen with navigation similar to drone driving. They can also fire fireballs at players by locking onto them on the mobile screen. Players often make dramatic movements to avoid being hit. The goal is to defeat the dragon by reducing its HP to zero. Each round lasts approximately 5 minutes — if time expires, the dragon wins; if the dragon’s HP reaches zero, the wizards win.

### v2

@@ 2.3. Co-located Mixed Reality
The focus of MR research has increasingly shifted from single-user experiences to multi-user collaborations, particularly within the CSCW community. Billinghurst and Kato coined the term “Collaborative Mixed Reality” to describe this emerging field (Billinghurst and Kato, 2002). Collaborative MR can be broadly categorized into remote collaboration and face-to-face co-located collaboration. Remote collaboration has been extensively studied (Schäfer et al., 2023). These studies have demonstrated MR’s potential to bridge physical distances and enhance communication in distributed teams. On the other hand, face-to-face co-located collaboration in MR (Billinghurst et al., 2002) has been explored in two main areas: interactions between MRHMD wearers (Nilsson et al., 2009) and collaborative work involving both handheld devices and MRHMDs (Zaman et al., 2023). The latter has particularly focused on bystander inclusion, addressing the challenge of integrating non-HMD users into MR experiences (Zhu and Grossman, 2020; Denning et al., 2014). Research on co-located collaboration between HMD wearers has investigated shared workspace designs, collaborative problem-solving, and social dynamics in MR environments (Sereno et al., 2020). Auda et al. (Auda et al., 2023) conducted a comprehensive scoping survey on these cross-reality systems.
Unlike traditional mobile games that rely on screen interactions—such as location-based multiplayer experiences like Pokémon GO (Xu et al., 2023; Paavilainen et al., 2017; Alha et al., 2019; Poretski et al., 2021) or co-located handheld AR games (Yusof et al., 2019; Reig et al., 2023; Xu et al., 2008; Nijholt, 2021)—MRHMD wearers naturally use their bodies as controllers. This shift from screen-touching to movement-based co-located bodily play represents a significant change in technology-mediated gameplay. While extensive research has explored co-located MR interactions (Bhattacharyya et al., 2019; Sereno et al., 2020; Dagan et al., 2022; Lin et al., 2024; Miedema et al., 2019; Grandi et al., 2019; Zhou et al., 2019; Radu et al., 2021; Thomas, 2012; He et al., 2019; D’Souza et al., 2018; Reig et al., 2023), these studies have primarily been conducted in controlled, private settings. Those studies highlight the need for designs that foster social inter-bodily engagement (Mueller et al., 2018a). Notably, commercial location-based entertainment like HADO (Araki et al., 2017), a multiplayer co-located dodgeball-like mixed reality sport, and Spatial Ops (Games, 2024), marketed as the world’s first competitive multiplayer shooter that “turns your home into a battlefield”, have emerged. However, these games require specific setups and predefined environments—such as Wi-Fi connectivity or special marker walls—which limit their potential for spontaneous play in public or outdoor settings. Despite some early academic explorations (Bonfert et al., 2017; Piekarski and Thomas, 2002), to our knowledge, few commercially available or open-source systems currently explore co-located bodily play using MRHMDs in public spaces.
2.4. Mixed Reality Concerns in Public Spaces
@@
In recent years, MRHMDs with passthrough functionality (Itoh et al., 2021) have sparked a notable shift from private, controlled settings to more public, spontaneous environments. These devices, including video and optical see-through technologies like Apple Vision Pro, Magic Leap, Microsoft HoloLens, and Meta Orion, enable users to view their physical surroundings while interacting with virtual elements (Radu et al., 2021; Bailenson et al., 2024). The rise in MRHMD usage in public settings raises social concerns for both wearers and non-wearers (Gugenheimer et al., 2017; Sapkota et al., 2021). The social acceptance (Koelle et al., 2020) of MRHMD usage in public spaces tends to decline significantly (Koutromanos and Kazakou, 2023; Profita et al., 2016; Sun and Schöning, 2025; Schwind et al., 2018; Eghbali et al., 2019; Gugenheimer et al., 2019, 2022; O’Hagan et al., 2023; Lebeck et al., 2018) due to three main factors:
@@
MRHMDs grant wearers capabilities—such as data overlays or video recording—that non-wearers cannot see or fully understand. This imbalance can negatively affect social experiences, privacy, and safety (Chung et al., 2023). Denning et al. (Denning et al., 2014) examined bystander privacy in AR, while O’Hagan et al. (O’Hagan et al., 2021, 2022) studied bystanders’ privacy concerns and awareness regarding AR glasses. A notable example of these social tensions is Google Glass, which faced criticism for prioritizing user capabilities over broader social impacts. This led to the term “Glasshole”, illustrating bystanders’ perception of antisocial behavior by Google Glass wearers (Due, 2015).
Embodied interactions
MRHMDs often rely on embodied interactions (Handosa et al., 2018), such as mid-air gestures and full-body movements. Although these interactions can be intuitive for manipulating virtual objects, they may also contravene established social norms, engendering social awkwardness (Plakias, 2024; Krauß et al., 2024; Chen et al., 2024; Guo et al., 2024). Moreover, these conspicuous gestures (Montero et al., 2010; Ahlström et al., 2014) may exacerbate the power imbalance between wearers and bystanders, as onlookers may feel excluded or uncertain about the wearer’s actions in mixed reality.
Social acceptance and safety in public spaces
When MRHMDs are used “on the go” in public spaces (Schwind et al., 2018), such as urban streets (Guo et al., 2024), public libraries (Kaeder et al., 2024), public transit (Bajorunaite et al., 2025, 2023; Medeiros et al., 2022; Bajorunaite et al., 2024; Medeiros et al., 2024a), and airplanes (Ng et al., 2021; Williamson et al., 2019), social acceptance, trust, security, and privacy become even more pressing (Katins et al., 2024; Millard et al., 2024; Roesner et al., 2014; Zhu et al., 2024b; Abraham et al., 2022). Interactions and movement in public traffic environments can pose additional risks, especially when MRHMD wearers navigate between physical and virtual elements. This interaction style intensifies the power imbalance and increases the potential for accidents or social friction with bystanders.
@@ 2.5. Research Gap: Immersive Mixed Reality Street Play
While the concept of MR street play has been advocated for several years, there remains a significant gap in research regarding co-located bodily play with MRHMDs in public spaces. To our knowledge, no commercial MRHMD games currently exist that can be reliably deployed or played on the streets. Although extensive empirical research has examined the social implications of smartphone-based street play like Pokémon GO (Alha et al., 2019), our literature review reveals no structured understanding or empirical studies of IMRSP. This under-exploration appears to stem from three main reasons:
Inaccessibility of affordable MRHMDs
As of 2025, MRHMDs with either full-color video passthrough or optical see-through—such as Vision Pro, HoloLens, and Meta’s Orion—remain prohibitively expensive, with prices starting at \$3,500 per device. This high cost poses a significant barrier to co-located multiplayer play, which requires multiple MRHMDs. Consequently, research in this area is typically confined to controlled lab settings, limiting the exploration of in-the-wild applications.
Limited technology for spontaneous, low‑latency multiplayer mixed reality co-location
Multiplayer mixed reality co-location requires low-latency spatial pose transmission—less than 20 ms latency—to prevent noticeable lag (Struye et al., 2020). This necessitates either deploying a Wi-Fi router in the wild or utilizing low-latency cellular networks like 5G, which do not yet provide ubiquitous coverage. Furthermore, routing data through remote servers between players often results in latency that exceeds user tolerance thresholds.
Social acceptance barriers and lack of established social norms
@@
2.6. Methods for Exploring Future Mixed Reality
This paper interrogates the prospective social implications of IMRSP as a form of play in public spaces in a future where MRHMDs are as ubiquitous as today’s smartphones. Currently, no commercially viable IMRSP experiences can be deployed on real streets with ecological validity. Therefore, we turn to design futuring (Fry, 2018) methods—approaches intended to provoke reflection on possible sociotechnical futures, reveal tensions and frictions (Pierce, 2021), challenge assumptions, and stimulate critical dialogue (Celik and Kaya, 2025).
Traditional design futuring (Fry, 2018; Kozubaev et al., 2020; Celik and Kaya, 2025; Lindley and Coulton, 2015) practices, including speculative design (Auger, 2013; Dunne, 2013; Barendregt and Vaage, 2021), design fiction (Sterling, 2009; Blythe, 2014), experiential futures (Candy and Dunagan, 2017), critical design (Bardzell et al., 2012), material speculation (Wakkary et al., 2015), typically employ fictional or low‑fidelity artifacts (Zhu et al., 2024a; Gaver, 2012), such as diegetic prototypes (Kirby, 2010), mock‑ups (Blythe et al., 2016), scripted scenarios (Candy and Dunagan, 2017), games (Coulton et al., 2016; Brown et al., 2016), or concept videos (Briggs et al., 2012; Odom et al., 2024; Sturdee et al., 2016) to provoke reflection, to imagine alternative worlds, and to foreground the discursive power. Such formats, however, often leave audiences on the far side of what Candy calls an “experiential gulf” (Candy, 2010): viewers discuss a hypothetical future rather than live it, producing debate rather than situated behavioral data (Ringfort-Felner et al., 2025). Elsden et al. (Elsden et al., 2017) further caution that lab‑based user enactments (Odom et al., 2012) can break the suspension of disbelief. To address this gap and facilitate experiential participation, Speculative Enactments (Elsden et al., 2017) stage scenarios as lived experiences that foreground real‑world consequences. Building on this, Immersive Speculative Enactments (Simeone et al., 2022) leverage virtual reality technologies to place participants in open‑ended situations that invite authentic action rather than following fixed linear narratives, as is common in Immersive Design Fiction (McVeigh-Schultz et al., 2018). Cools et al. (Cools et al., 2024) offer an Immersive Speculative Enactment example in which participants experience fictional MR contact lens scenarios through a multi-user VR simulation, generating experiential insights into the technology’s potential impact on daily life.
Building on this lineage, we treat the future not as a scenario to be imagined but as a situation to be enacted and studied in situ. Unlike many design futuring approaches, IMRSP is not a speculative impossibility; it is technically achievable today. We see opportunities to directly implement IMRSP through current capabilities of hardware and software systems. What limits our ability to observe its social implications is not feasibility but adoption: MRHMDs are not yet ubiquitous in everyday public life. We could implement high‑fidelity, fully functional IMRSP prototypes to bring the future into the present and use them as research probes in public spaces, allowing us to study its real-world implications directly.
This performative, embodied approach (Buchenau and Suri, 2000; Benford et al., 2013) turned city streets into staged yet plausible near‑future environments in which passersby could actively participate. Through in‑the‑wild trials (Rogers and Marshall, 2017), we documented real‑time social behaviors, safety negotiations, and bystander–player dynamics—insights that fictional formats cannot surface. Our study thus transforms a what‑if into a lived experiment: a form of speculative enactment in which participants “act on and experience the speculation, beyond commenting on it(Elsden et al., 2017). Conducted as Research through Design (RtD) (Zimmerman et al., 2007), the field deployment yields ecologically valid, embodied behavioral data and shows how combining speculative futures with rigorous in situ experimentation can both reveal and shape trajectories for emerging technologies.
@@ 3. System Design
To explore the real-world social implications of IMRSP, we employed a research-through-design approach. We developed a series of game prototypes called “Multiplayer Omnipresent Fighting Arena” (MOFA) and deployed these games in real-world scenarios. MOFA serves as an open-source MR game framework that enables co-located, multiplayer bodily play through MRHMDs, allowing spontaneous play on the street via peer-to-peer local wireless networking.
@@ 3.1.1. Easy to Understand and Learn
The game probe should be easily understood by city dwellers on the streets. Inspired by the Harry Potter fantasy, particularly wizard dueling, we adopted a ‘magic’ theme for the gameplay. This aligns well with the characteristics of augmented reality media—unseen energy invisible to “Muggles”, who are unaware of magic. This theme draws from popular culture that has been deeply rooted in global consciousness for over three decades. It makes the rules and settings intuitively understandable for many without explanation in “in-the-wild” settings. Even when behaviors might seem unusual in public, they are explainable and comprehensible because people have been familiarized through Harry Potter movies—they recognize these wand dueling interactions. Players can instantly immerse themselves in imaginative wizard and witch roles.
@@ 3.1.2. Physical Props for Social Affordance
All gameplay for MRHMD wearers in MOFA incorporates a physical prop: a wand. We discovered that the wand provides an intuitive affordance for players to understand the hand-waving gesture used to cast spells. This proves effective for three primary reasons: (1) It aligns with the stereotypical image of wizardry popularized by the Harry Potter series, enabling people to easily grasp the interaction without tutorials; (2) The physical shape and weight of the wand facilitate a natural “wave down” gesture, which is more intuitive than any bare-handed gesture for casting spells (which typically require tutorials). With wands, the interaction becomes second nature; (3) The wands help passersby readily comprehend the context of the games, even without seeing the augmented reality imaging overlay. Addressing the social acceptance concerns described in Section 2.4, the wand serves as a strong social affordance (Isbister et al., 2018) for both players and passersby in public spaces.
3.1.3. Co-located Bodily Play without Physical Contact
@@
As envisioned in promotional videos, MRHMD wearers engage in embodied interactions involving full-body movements, transcending traditional screen-touching gameplay. MOFA exemplifies co-located bodily play with MRHMDs, showcasing interactions ranging from single-player to multiplayer scenarios. The game deliberately avoids physical contact for two key reasons: (1) many existing mixed reality systems cannot detect physical contact between players, making tangible interactions like melee attacks impractical in game design; (2) to prioritize player safety, we designed the game with ranged attacks only, eliminating the risk of direct physical contact between players. MOFA strategically incorporates a magic shield in front of players to indicate an attacking target, which allows opponents to aim at the shield, providing a clear target without risking physical harm. Unlike traditional fighting games, where avatars display damage through fainting, blood spatter, or falling, MOFA uses a shattering shield visual effect to provide feedback when an opponent successfully lands a hit. This design ensures both safety and clear interaction feedback.
3.1.4. Omnipresent Co-location
@@
MOFA is designed for deployment anywhere, transforming any street into a game arena—hence its name, “omnipresent”. It should operate without Wi-Fi or cellular connections. Mixed reality co-location requires low-latency spatial pose transmission (under 20 ms) to prevent noticeable lag. To address the research gap in spontaneous co-location technology outlined in Section 2.5, we employed an open-source co-location toolkit (Hu et al., 2023a). This toolkit utilizes Multipeer Connectivity [^6] https://developer.apple.com/documentation/multipeerconnectivity, accessed on October 25, 2024, the technology powering Apple’s AirDrop, to enable local networking. This innovation allows street play to begin anywhere, even in rural areas lacking cellular signals or Wi-Fi access. The technology also features a rapid QR scanning process for synchronizing coordinates across devices, eliminating the need to pre-scan the physical environment and enabling spontaneous player as well as spectator participation.
3.2. Gameplay Overview
Figure 2. Detailed game and user experience design, using “The Duel” as an example. It illustrates: (1) the first-person perspective of the game with all UI elements users can view through the MR interface; (2) interaction design details using Apple Watch to detect spell-casting gestures; (3) magic spell and shield system design; (4) co-location and networking user flow: how two players connect and how spectators join to watch in third-person view; (5) detailed round-based dueling game mechanism design; and (6) conversion process from monocular screen-based AR view into immersive stereoscopic MR view based on HoloKit.
MOFA includes three co-located bodily game modes. Gameplay specifications are provided in Table 1. Detailed game and user experience design are depicted in Fig. 2.
Table 1. Specifications of three MOFA game modes
@@ 3.2. Gameplay Overview
Required Number of Players 1 2 3 or more (with one player controlling the Dragon via a mobile device)
Duration per Round 90 seconds 90 seconds Approximately 5 minutes
@@
3.2.1. “The Training” - a single-player gameplay mode
@@
“The Training” is a single-player dueling mode where players engage in magical battles with an AI-controlled cartoon opponent (see Fig. 1a). Inspired by Harry Potter’s training scenes [^7] https://harrypotter.fandom.com/wiki/Cast-a-Spell_training_room, visited on October 25, 2024, players take on the role of wizards, honing their wand skills to cast spells and attack virtual ghosts in Hogwarts. This game probe focuses on one spell-casting interaction: detecting “waving down” hand gestures with physical wands. It is a single-player MRHMD gameplay in public spaces.
Game flow of “The Training”
@@
First, the player puts on the Apple Watch while holding a wand and launches the game app on both the watch and iPhone. After selecting “The Training” mode, the player inserts the iPhone into the headset. Once wearing the headset and detecting a floor plane, the player taps the Apple Watch to place the virtual AI cartoon avatar opponent on the floor in front of them. Upon the audio cue “1, 2, 3, start!” the round begins, and the AI starts casting spells and dodging. The player dodges incoming spells and counterattacks by waving their watch-worn hand, which triggers spells through motion detection. Players earn one point by completely breaking their opponent’s mixed reality shield. The player tries to break the AI opponent’s shield as many times as possible within 90 seconds; the player wins if their score exceeds the AI’s.
3.2.2. “The Duel” - a competitive gameplay mode
@@
“The Duel” is a competitive gameplay mode pitting two players against each other in magical combat. It expands on “The Training” mechanics, transitioning from player-versus-AI to player-versus-player (see Fig. 1b). Drawing inspiration from Harry Potter’s classic Wizard Dueling [^8] https://harrypotter.fandom.com/wiki/Duelling, visited on October 25, 2024, players use wands to cast spells and fight each other, incorporating dodgeball-like body movements to avoid attacks. This simulates future scenarios where MRHMD users might encounter each other on the street for impromptu duels.
Game flow of “The Duel”
@@
First, two human players launch the game app on both their watches and iPhones. The host player selects “The Duel” mode. A QR code is displayed on the phone. The client player scans the host’s QR code with their iPhone. After confirming the connection, both players insert their iPhones into their headsets. Once wearing the headsets, the two players face each other and click their Apple Watches to start the game round. Upon the audio cue “1, 2, 3, start!” the players cast spells, dodge, and counterattack. Players earn one point by completely breaking their opponent’s mixed reality shield. Each round lasts 90 seconds, and the player with the most shield breaks wins. After each round, players can check their Apple Watch to see their score, hit rate, estimated distance moved, and estimated calories burned. We observed that players often make dramatic movements to avoid being hit. Thus, we selected 90-second rounds to create a mild exertion game suitable for players with average fitness levels, burning approximately 10 kilocalories per round (see Fig. 3e).
Figure 3. Top: Device-pairing diagram for MOFA “The Dragon”. The image on the right shows a frame from a spectator video recorded on Device 6. Bottom: Step-by-step onboarding process: (a) All participants launch the game app on their iPhones. (b) The host device initiates the game session, and guest devices join by scanning a QR code displayed on the host’s screen. (c) Wizard players switch to stereoscopic mixed reality mode and insert their phones into the headsets. (d) Wizard players press the “Ready” button on their watches to activate their powers and enter the game. (e) After each round, performance data—“Score” (kills), “Hit Rate”, “Distance” (distance moved), and “Calorie” (calories burned)—are displayed on each wizard player’s Apple Watch. Players can tap “Play Again” on the watch to start a new round.
3.2.3. “The Dragon” - a collaborative and puppeteering gameplay mode
@@
“The Dragon,” designed for three or more players and extending from “The Duel”, introduces a new interaction paradigm (see Fig. 1c): puppeteering. In this mode, one handheld AR player—dubbed the “puppeteer”—controls a virtual dragon using their mobile phone, interacting with MRHMD users. Meanwhile, players equipped with MRHMDs become dragon-hunting warriors, casting spells to battle the dragon while using body movements to avoid attacks. This gameplay is inspired by Game of Thrones’ famous dragon‑hunting scenes [^9] https://game-of-thrones-winter-is-coming-game.fandom.com/wiki/Dragon_Hunt, visited on October 25, 2024. It envisions a transitional future where MRHMD users coexist with handheld AR users, enabling collaborative play between different device types when they encounter each other on the street.
Game flow of “The Dragon”
@@
The host player puppeteers the dragon on their mobile screen (Device 1 in Fig. 3). First, the host selects “The Dragon” mode and displays a QR code on their phone screen. Three human players (Devices 2, 3, 4 in Fig. 3) scan the host’s QR code with their iPhones (see Fig. 3a-b). After confirming the connection, all players insert their iPhones into their headsets (see Fig. 3c). Two spectators (Device 5 and 6 in Fig. 3) may also scan the QR code with their iPhones to join as spectators. Once wearing the headsets, the three players face the designated play area after confirming with the host player and tap “Ready” on their Apple Watches to indicate readiness (see Fig. 3d). The host uses their phone to detect the floor plane, then spawns the virtual dragon to begin the battle. Upon the audio cue “1, 2, 3, start!”, the round begins. Players cast spells to attack the dragon while dodging its counterattacks. The host controls the dragon via touchscreen input on their phone, using navigation similar to that of drone driving. They can also launch fireballs at players by locking onto them on the mobile screen. Similar to the gameplay of “The Duel”, players often make dramatic movements to avoid being hit. The goal is to defeat the dragon by reducing its health points (HP) to zero. Each round lasts approximately 5 minutes — if time expires, the dragon wins; if the dragon’s HP reaches zero, the wizards win.